OpenStudy (emmigrace222):

\(match~the~answers\) 42 + 3 - 2(3) A) 18 22 • 8 ÷ 2 + 2 B) 13 (42 + 16) + 12 C) 44 8 • 32 + 3 D) 75

OpenStudy (emmigrace222):

@undeadknight26

TheSmartOne (thesmartone):

42+3-2(3)=42+3-6=39 Which is not a choice..

OpenStudy (anonymous):

42 + 3 - 2(3) should equal 39 22 • 8 ÷ 2 + 2 = 90 (42 + 16) + 12 = 70 8 • 32 + 3 = 259 Neither match

TheSmartOne (thesmartone):

22x8=176 176/2= 88 88+2=90 Which is also not a choice

TheSmartOne (thesmartone):

None of your answer choices match for some reason...

OpenStudy (emmigrace222):

ugh

OpenStudy (emmigrace222):

thats answer i have to use

OpenStudy (emmigrace222):

@undeadknight26

undeadknight26 (undeadknight26):

Dey dont match! *throws computer at wall* WHYYYYY

OpenStudy (emmigrace222):

idk!! D:

OpenStudy (emmigrace222):

UGH!!!!! FML!!!

undeadknight26 (undeadknight26):

22 • 8 ÷ 2 + 2 = 44 Only one that matches.

OpenStudy (emmigrace222):

urgh

OpenStudy (anonymous):

44 is not correct. You add at the end

OpenStudy (emmigrace222):

-flips table- urgh